19. Calculate the client’s fluid maintenance? His weight is 18 kg.

The hourly fluid requirement: __________/ hour; The daily fluid requirement: ___________/day;

(4ml/kg for the first 10kg) + (2ml/kg for the second 10 kg) + (1ml/kg for every kg above 20) = hourly rate

(4ml x 10kg) + (2ml x 10kg) + (1ml x 5kg) = hourly rate

40ml + 20ml + 5ml = 65ml/hr Using the 4/2/1 method, this child's hourly maintenance fluid rate is 65mL/hr

20. Aspirin is ordered for your patient. The safe therapeutic range is 80 mg – 100 mg/day/day divided into 4 doses until fever resolves. a.) Is the dose within therapeutic range?

b.) What is the rationale for the ‘high dose’ range for this child?

c.) On hand is Aspirin 81 mg/tablet. How many tablets does the child receive? ___________ tablets
Answer #1

The weight of the client=18 kg

hourly fluid requirement =

(as described in the formula ,ml/hour , 4 ml/kg till 10 kg.

2 ml/kg for the next 10 kg

1ml/kg remaining weight above 20 kg

thus,(4 ml* 10kg)+(2ml* 8kg)=40+16= 56ml/hour

hourly fluid requirement=56 ml/hour

therefore, daily fluid requirement= (if it is 56 for 1 hour , then for 24 hour=

= 56 ml*24 = 1344 ml/hour